OpenStudy (anonymous):

tanx(1-sin^2x)=1/2sin2x

OpenStudy (anonymous):

\[1-\sin ^x=\cos ^2x,\tan x=\frac{ \sin x }{ \cos x },\sin 2 x=2\sin x~\cos x\]